OpenStudy (anonymous):

the line through A (6,6) parallel to the y -axis intersects the line m: 2x-y=-3 at B. Find the coordinates of B

OpenStudy (anonymous):

the coordinates should be (6 ,15)

OpenStudy (anonymous):

since the line is parallel to the y axis, its equation will only pass through x = 6. So the equation of the first line should be x =6. It intersects a second line 2x-y = -3. Substitute x = 6 into the second equation and you get 12 - y = -3 -y = -3 -12 -y = -15 y = 15 so the coordinates are (6 , 15)

OpenStudy (lgbasallote):

if you were going to give a solution...why didnt you just give it? why say the answer first then give a solution? i just see no sense in it...answer before solution...